linda reads 1/5 of a book on friday and 2/7 of the book on saturday. On sunday she read 5/9 of the remaining pages. What fraction of the book did she read on sunday?
I have tried to answer this question multiple times, but I can't seem to get it. And its due tomorrow! Please help. Thank you.

Answer by rfer(16322)

1/5+2/7=
7/35+10/35=17/35
35/35-17/35=18/35
------------------
18/35*5/9=90/315=2/7 read Sunday